### v1

@@ Abstract
Inexact proximal augmented Lagrangian methods (pALMs) are particularly appealing for tackling convex constrained optimization problems because of their elegant convergence properties and strong practical performance. To solve the associated pALM subproblems, efficient methods such as Newton-type methods are essential. Consequently, the effectiveness of the inexact pALM hinges on the error criteria used to control the inexactness when solving these subproblems. However, existing inexact pALMs either rely on absolute-type error criteria (which may complicate implementation by necessitating the pre-specification of an infinite sequence of error tolerance parameters) or require an additional correction step when using relative error criteria (which can potentially slow down the convergence of the pALM). To address this deficiency, this paper proposes ripALM, a relative-type inexact pALM, which can simplify practical implementation while preserving the appealing convergence properties of the classical absolute-type inexact pALM. We emphasize that ripALM is the first relative-type inexact version of the vanilla pALM with provable convergence guarantees. Numerical experiments on quadratically regularized optimal transport (OT) problems demonstrate the competitive efficiency of the proposed method compared to existing methods. As our analysis can be extended to a more general convex constrained problem setting, including other regularized OT problems, the proposed ripALM may provide broad applicability and has the potential to serve as a basic optimization tool.
Keywords: augmented Lagrangian method; proximal term; relative-type error criterion; asymptotically Q-(super)linear convergence rate; quadratically regularized optimal transport
@@ 1 Introduction
where $f:\mathbb{R}^{N}\to\mathbb{R}\cup\{+\infty\}$ is a (possibly nonsmooth) proper closed convex function, $A\in\mathbb{R}^{M\times N}$ and $\bm{b}\in\mathbb{R}^{M}$ are given data. While we focus on (1.1) in this paper, we would like to mention that the inexact algorithmic framework along with the associated theory developed in this paper can be extended to a more general convex constrained optimization problem studied in [11].
@@
which consists of minimizing one ALM-subproblem to get $\bm{x}^{k+1}$ and one gradient ascent step (associated with $\mathcal{L}_{\sigma_{k}}^{\rm prim}$) with step size $\sigma_{k}$ to get $\bm{y}^{k+1}$. Since the augmented Lagrangian function $\mathcal{L}_{\sigma}^{\rm prim}(\cdot)$ is associated with the primal problem (1.1), the above ALM is usually named as a primal-based ALM. Alternatively, one may also consider a dual-based ALM which is associated with the dual problem of (1.1) given by (modulo a minus sign)
@@
where $\mathtt{M}_{\sigma f}(\bm{x}):=\min\limits_{\bm{y}}\left\{f(\bm{y})+\frac{1}{2\sigma}\|\bm{y}-\bm{x}\|^{2}\right\}$ denotes the Moreau envelope of the function $\sigma f(\cdot)$ at $\bm{x}$. Consequently, for a given sequence of penalty parameters $\{\sigma_{k}\}\subseteq\mathbb{R}_{++}$ and an initial point $\bm{x}^{0}\in\mathbb{R}^{N}$, the dual-based ALM reads as follows:
$\begin{cases}\bm{y}^{k+1}\in\arg\min\limits_{\bm{y}\in\mathbb{R}^{M}}\left\{\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y},\bm{x}^{k})\right\},\\[5.0pt] \bm{x}^{k+1}=\bm{x}^{k}+\sigma_{k}\nabla_{\bm{x}}\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y}^{k+1},\bm{x}^{k})=\texttt{prox}_{\sigma_{k}f}\big{(}\bm{x}^{k}+\sigma_{k}{A}^{\top}\bm{y}^{k+1}\big{)}.\end{cases}$
@@
The applicability of the ALM has been significantly expanded since Rockafellar’s seminal works [38, 39], which established a deep connection between the ALM and the proximal point algorithm (PPA). Moreover, the elegant convergence properties of the ALM can be established under the Lipschitz continuity of a certain solution mapping. Subsequent works have shown that this stringent condition can be further relaxed; see, for example, [8, 29] and references therein. Indeed, understanding the convergence properties of the ALM continues to be a topic of interest in the literature. Apart from theoretical advancements, numerous studies have demonstrated the remarkable efficiency of the ALM in solving various convex problems, including conic programming problems [20, 22, 24, 55], statistical optimization problems [19, 25, 53], optimization problems in machine learning [47, 48] and image/signal processing [18, 27], to mention just a few. These successful applications highlight the critical roles of both the solution methods used for solving ALM-subproblems and the error criteria that control the accuracy required for solving these subproblems, while preserving the appealing convergence properties of the ALM.
@@
$\begin{cases}\bm{y}^{k+1}=\arg\min\limits_{\bm{y}\in\mathbb{R}^{M}}\Big{\{}\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y},\bm{x}^{k})+\cfrac{\tau_{k}}{2\sigma_{k}}\left\|\bm{y}-\bm{y}^{k}\right\|^{2}\Big{\}},\\[8.0pt] \bm{x}^{k+1}=\bm{x}^{k}+\sigma_{k}\nabla_{\bm{x}}\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y}^{k+1},\bm{x}^{k})=\texttt{prox}_{\sigma_{k}f}\big{(}\bm{x}^{k}+\sigma_{k}{A}^{\top}\bm{y}^{k+1}\big{)}.\end{cases}$ (1.3)
@@
To make the pALM truly implementable and practical, it must allow approximate solutions to the subproblem with progressively improved accuracy, and the associated error criterion must be practically verifiable, while preserving desirable convergence properties of the outer loop. To achieve this, an absolute-type error criterion was introduced by Rockafellar for the pALM in [38] and is now widely used in the literature; see, for example, [21, 23, 25, 26, 33]. However, this absolute-type error criterion requires the specification of a summable sequence of tolerance parameters, which must be tuned to avoid being too conservative or too aggressive. As a result, careful parameter tuning of this sequence is often necessary to achieve superior convergence performance. This process can result in excessive effort and potential inefficiencies in practical implementations. A notable alternative is to adopt a relative-type error criterion, which seeks to control the errors in minimizing the subproblems based on some quantity related to the progress of the algorithm. This approach eliminates the need for a summable tolerance sequence, thereby enhancing practical implementability. The adoption of a relative-type error criterion started from the seminal works of Solodov and Svaiter [44, 42, 43, 45, 46] on inexact versions of the PPA, and has since influenced the development of inexact versions of numerous algorithms, including the variable metric PPA [30], ALM [11], ADMM [12], FISTA [3], and the Bregman-type methods [49, 52]. However, to the best of our knowledge, research on relative-type inexact versions of the pALM remains in its early stage, with limited exploration in [10, 17, 51]. In particular, all these inexact versions are derived by applying a (variable metric) hybrid proximal extragradient method [30, 42, 43] to a certain primal-dual solution mapping of the convex constrained problem, and therefore require an additional correction step to ensure convergence. Consequently, the algorithms developed in [10, 17, 51] essentially deviate from the vanilla pALM framework (1.3). Moreover, our numerical results also indicate that the correction step tends to slow down the convergence of the pALM and thus increase the computational cost. These observations naturally lead us to raise the following question:
@@
In light of this, the primary goal of this paper is to address the above question by designing a relative-type error criterion for the vanilla pALM (1.3). Note that the relative-type error criterion proposed by Eckstein and Silva [11] for the ALM innovatively introduces an auxiliary error variable to control the inexactness while ensuring convergence. In this work, we shall demonstrate that a similar idea can be adapted to the vanilla pALM, but with essential modifications. We should emphasize that the corresponding theoretical convergence analysis also requires substantial modifications, due to the inclusion of the proximal term. Interestingly, while the initial motivation for considering pALM was to ensure the nonsingularity of the (generalized) Hessian of the ALM subproblem’s objective function at its solution point, our analysis reveals that this proximal term can further improve the existing convergence properties of the relative-type inexact ALM studied in [1, 11, 54]. Specifically, under some common assumptions, it ensures the convergence of both the primal and dual sequences, as well as their asymptotically Q-(super)linear convergence rates. The key contributions and findings of this paper are summarized as follows:
@@
  • We develop a relative-type inexact proximal augmented Lagrangian method (ripALM) to solve the dual problem (1.2). This proposed ripALM is the first inexact version of the vanilla pALM (1.3). It not only shares the same convergence properties for the sequence $\{\bm{x}^{k}\}$ as the relative-type inexact ALM studied in [1, 11, 54], but also offers other theoretical advantages in that the sequence $\{\bm{y}^{k}\}$ is guaranteed to converge, and both primal and dual sequences achieve asymptotically Q-(super)linear convergence rates under a relatively weaker error bound assumption. These results provide strong theoretical support for the applicability of ripALM, particularly in scenarios where accurate solutions of both primal and dual problems are required.
@@
  • Numerically, we implement the proposed ripALM and apply it to solve quadratically regularized optimal transport (QROT) problems. Extensive numerical results validate the promising performance of the proposed ripALM for solving large-scale QROT problems. Moreover, comparisons with existing methods underscore our motivation and contributions in developing a relative-type error criterion for the vanilla pALM. We also emphasize that our ripALM can be extended to solve optimal transport problems with other convex regularizers such as the group quadratic regularizer [51], provided that the regularizers are proximal-friendly and the generalized Jacobians associated with the pALM-subproblems are easy to compute. Given the growing interests in developing efficient algorithmic frameworks for computational optimal transport [31], our work may open up new possibilities for applying the powerful pALM algorithmic framework in a wide variety of real-world applications.
@@
The remaining parts of this paper are organized as follows. Section 2 describes the main algorithmic framework of ripALM, whose convergence analysis is conducted in Section 3. Section 4 showcases how to apply the proposed ripALM for solving QROT problems. Numerical experiments are conducted in Section 5. Finally, some concluding remarks are summarized in Section 6.
@@
Notation. We use $\mathbb{R}^{n}$, $\mathbb{R}_{+}^{n}$, $\mathbb{R}^{m\times n}$ and $\mathbb{R}_{+}^{m\times n}$ to denote the sets of $n$-dimensional real vectors, $n$-dimensional non-negative vectors, $m\times n$ real matrices and $m\times n$ real non-negative matrices, respectively. For a vector $\bm{x}\in\mathbb{R}^{n}$, $x_{i}$ denotes its $i$-th entry, $\|\bm{x}\|$ denotes its Euclidean norm, and $\|\bm{x}\|_{H}:=\sqrt{\langle\bm{x},\,H\bm{x}\rangle}$ denotes its weighted norm associated with a symmetric positive definite matrix $H\in\mathbb{R}^{n\times n}$. For a matrix $X\in\mathbb{R}^{m\times n}$, ${X_{ij}}$ denotes its $(i,j)$-th entry, $\|X\|_{F}$ denotes its Frobenius norm, and $\texttt{vec}(X)$ denotes the vectorization of $X$, where $[\texttt{vec}(X)]_{i+(j-1)m}=X_{ij}$ for any $1\leq i\leq m$ and $1\leq j\leq n$. For simplicity, given an integer $n>0$, we use $\bm{1}_{n}\in\mathbb{R}^{n}$ to denote the $n$-dimensional vector of all ones, and use $I_{n}$ to denote the $n\times n$ identity matrix.
@@ 2 A relative-type inexact pALM
In this section, we focus on developing a relative-type inexact proximal augmented Lagrangian method (ripALM) to solve the dual problem (1.2). The algorithmic framework is developed based on the parametric convex duality framework (see, for example, [36, 37] and [40, Chapter 11]).
@@
Note that $G$ is proper closed convex since $f^{*}$ is proper closed convex. Then, the (ordinary) Lagrangian function of problem (1.2) can be defined by taking the concave conjugate of $G$ with respect to its second argument (see [40, Definition 11.45]), that is,
@@
Clearly, $\ell$ is convex in its first argument and concave in the second argument. For a given penalty parameter $\sigma>0$, the augmented Lagrangian function of problem (1.2) is defined by (see [40, Example 11.57])
$\mathcal{L}_{\sigma}(\bm{y},\,\bm{x}):=$ $\sup\limits_{\bm{s}\in\mathbb{R}^{N}}\left\{\ell(\bm{y},\bm{s})-\frac{1}{2\sigma}\|\bm{s}-\bm{x}\|^{2}\right\}$
@@
With the above preparations, we are ready to present our ripALM for solving problem (1.2) in Algorithm 1.
@@
Input: $\rho\in[0,1)$, $\{\sigma_{k}\}_{k=0}^{\infty}$ is a sequence of real numbers such that $\inf_{k\geq 0}\{\sigma_{k}\}>0$, and $\{\tau_{k}\}_{k=0}^{\infty}$ is a sequence of real numbers such that $\inf_{k\geq 0}\{\tau_{k}\}>0$ and $\sup_{k\geq 0}\{\tau_{k}\}<\infty$. Choose $\bm{y}^{0},\,\bm{w}^{0}\in\mathbb{R}^{M}$ and $\bm{x}^{0}\in\mathbb{R}^{N}$ arbitrarily. Set $k=0$.
@@
    $\leq$ $\;\rho\left(\big{\|}\texttt{prox}_{\sigma_{k}f}\big{(}\bm{x}^{k}+\sigma_{k}{A}^{\top}\bm{y}^{k+1}\big{)}-\bm{x}^{k}\big{\|}^{2}+\tau_{k}\big{\|}\bm{y}^{k+1}-\bm{y}^{k}\big{\|}^{2}\right),$
@@
    $\bm{x}^{k+1}=\texttt{prox}_{\sigma_{k}f}\big{(}\bm{x}^{k}+\sigma_{k}{A}^{\top}\bm{y}^{k+1}\big{)},\quad\bm{w}^{k+1}=\bm{w}^{k}-\sigma_{k}\Delta^{k+1}.$ (2.6)
@@
In line with pALM-type methods, at each iteration, our ripALM in Algorithm 1 approximately minimizes the sum of the augmented Lagrangian function $\mathcal{L}_{\sigma_{k}}(\cdot,\,\bm{x}^{k})$ and a proximal term $\frac{\tau_{k}}{2\sigma_{k}}\big{\|}\cdot-\bm{y}^{k}\big{\|}^{2}$ under the error criterion (2.5), followed by the updates of the multiplier $\bm{x}^{k}$ and the auxiliary error variable $\bm{w}^{k}$. Due to the strong convexity of the objective function in (2.4), the subproblem (2.4) has a unique solution, denoted as $\bm{y}^{k,*}$, satisfying $\Delta^{k,*}:=\nabla_{y}\mathcal{L}_{\sigma_{k}}(\bm{y}^{k,*},\bm{x}^{k})+\tau_{k}\sigma_{k}^{-1}(\bm{y}^{k,*}-\bm{y}^{k})=\bm{0}$. Then, for any minimizing sequence $\{\bm{y}^{k,t}\}$ for (2.4) converging to $\bm{y}^{k,*}$, we have that $\Delta^{k,t}:=\nabla_{y}\mathcal{L}_{\sigma_{k}}(\bm{y}^{k,t},\bm{x}^{k})+\tau_{k}\sigma_{k}^{-1}(\bm{y}^{k,t}-\bm{y}^{k})\to\bm{0}$ and consequently $2\big{|}\langle\bm{w}^{k}-\bm{y}^{k,t},\,\sigma_{k}\Delta^{k,t}\rangle\big{|}+\big{\|}\sigma_{k}\Delta^{k,t}\big{\|}^{2}\to 0$ as $t\to\infty$. Meanwhile, if $\bm{y}^{k}$ is not the solution of (2.4) and $\tau_{k}\neq 0$, the right-hand side of (2.5) cannot be zero. Therefore, in this case, our error criterion (2.5) must hold after finitely many $t$ iterations, and thus it is achievable.
@@
Compared to the recent semismooth Newton based inexact proximal augmented Lagrangian (Snipal) method developed in [21, Section 3], our ripALM in Algorithm 1 employs a significantly different error criterion (2.5) for solving the subproblem (2.4). Specifically, in our context, Snipal requires the error term $\Delta^{k+1}$ to satisfy
[…]

### v2

@@ Abstract
Inexact proximal augmented Lagrangian methods (pALMs) are particularly appealing for tackling convex constrained optimization problems because of their elegant convergence properties and strong practical performance. To solve the associated pALM subproblems, efficient methods such as Newton-type methods are essential. Consequently, the effectiveness of the inexact pALM hinges on the error criteria used to control the inexactness when solving these subproblems. However, existing inexact pALMs either rely on absolute-type error criteria (which may complicate implementation by necessitating the pre-specification of an infinite sequence of error tolerance parameters) or require an additional correction step to guarantee convergence when using relative-type error criteria (which can potentially degrade the practical performance of pALM). To address these deficiencies, this paper proposes ripALM, a relative-type inexact pALM, for linearly constrained convex optimization. This method can simplify practical implementation while preserving the appealing convergence properties of the classical absolute-type inexact pALM. To the best of our knowledge, ripALM is the first relative-type inexact version of the vanilla pALM with provable convergence guarantees. Numerical experiments on quadratically regularized optimal transport problems and basis pursuit denoising problems demonstrate the competitive efficiency of the proposed method compared to existing state-of-the-art methods. As our model encompasses a wide range of application problems, the proposed ripALM offers broad applicability and has the potential to serve as a basic optimization toolbox.
Keywords: augmented Lagrangian method; proximal term; relative-type error criterion; asymptotic (super)linear convergence rate; quadratically regularized optimal transport; basis pursuit denoising
@@ 1 Introduction
where $f:\mathbb{R}^{N}\to\mathbb{R}\cup\{+\infty\}$ is a (possibly nonsmooth) proper closed convex function, $A\in\mathbb{R}^{M\times N}$ and $\bm{b}\in\mathbb{R}^{M}$ are given data. Problem (1.1) is a classical and widely studied template in convex optimization. Moreover, since the smoothness of the objective function $f$ is not required in this work, problem (1.1) actually encompasses a much broader class of application problems than it may appear at first glance. For instance, by leveraging an indicator function of a closed convex set, problem (1.1) naturally accommodates more general convex constraints; see also two illustrative examples in Section 4.
@@
which consists of minimizing one ALM-subproblem to get $\bm{x}^{k+1}$ and one gradient ascent step (associated with $\mathcal{L}_{\sigma_{k}}^{\rm prim}$) with a step size $\sigma_{k}$ to get $\bm{y}^{k+1}$. Since the augmented Lagrangian function $\mathcal{L}_{\sigma}^{\rm prim}$ is associated with the primal problem (1.1), the above ALM is usually named as a primal-based ALM. Alternatively, one may also consider a dual-based ALM which is associated with the dual problem of (1.1) given by (modulo a minus sign)
@@
where $\mathtt{M}_{\sigma f}(\bm{x}):=\min\limits_{\bm{z}}\left\{f(\bm{z})+\frac{1}{2\sigma}\|\bm{z}-\bm{x}\|^{2}\right\}$ denotes the Moreau envelope of the function $\sigma f$ at $\bm{x}$. Consequently, for a given sequence of penalty parameters $\{\sigma_{k}\}\subseteq\mathbb{R}_{++}$ and an initial point $\bm{x}^{0}\in\mathbb{R}^{N}$, the dual-based ALM reads as follows:
$\begin{cases}\bm{y}^{k+1}\in\arg\min\limits_{\bm{y}\in\mathbb{R}^{M}}\left\{\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y},\bm{x}^{k})\right\},\\[5.0pt] \bm{x}^{k+1}=\bm{x}^{k}+\sigma_{k}\nabla_{\bm{x}}\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y}^{k+1},\bm{x}^{k})=\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}{A}^{\top}\bm{y}^{k+1}\big).\end{cases}$ (1.3)
@@
The applicability of the ALM has been significantly expanded since Rockafellar’s seminal works [41, 42], which established a deep connection between the ALM and the proximal point algorithm (PPA). Moreover, the asymptotic (super)linear convergence rate of the ALM can be established under the assumption that the inverse of the subdifferential of the primal or dual objective function is Lipschitz continuous at the origin. Subsequent works have shown that this stringent condition can be further relaxed; see, for example, [11, 32] and references therein. Indeed, understanding the convergence properties of the ALM continues to be a topic of interest in the literature. Apart from theoretical advancements, numerous studies have demonstrated the remarkable efficiency of the ALM in solving various convex problems, including conic programming problems [23, 25, 27, 62], statistical optimization problems [22, 28, 60], optimization problems in machine learning [50, 54], and image/signal processing [21, 30], to mention just a few. These successful applications highlight the crucial roles of both the subsolvers employed to handle ALM subproblems and the error criteria used to control the accuracy required for solving these subproblems, while maintaining the desirable convergence properties of the ALM.
@@
$\begin{cases}\bm{y}^{k+1}=\arg\min\limits_{\bm{y}\in\mathbb{R}^{M}}\Big\{\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y},\bm{x}^{k})+\cfrac{\tau_{k}}{2\sigma_{k}}\left\|\bm{y}-\bm{y}^{k}\right\|^{2}\Big\},\\[8.0pt] \bm{x}^{k+1}=\bm{x}^{k}+\sigma_{k}\nabla_{\bm{x}}\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y}^{k+1},\bm{x}^{k})=\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}{A}^{\top}\bm{y}^{k+1}\big).\end{cases}$ (1.4)
@@
To make the pALM truly implementable and practical, it must allow approximate solutions to the subproblem with progressively improved accuracy, and the associated error criterion must be practically verifiable, while preserving desirable convergence properties of the outer loop. To achieve this, an absolute-type error criterion was introduced by Rockafellar for the pALM in [41] and is now widely used in the literature; see, for example, [24, 26, 28, 29, 36]. The term “absolute” means that the sequence of error tolerances, which controls the accuracy required for solving each pALM subproblem, is predetermined and remains fixed throughout the iterations. However, such a sequence typically needs to be summable and must be carefully tuned to avoid being either overly conservative or excessively aggressive. Consequently, achieving desirable numerical performance often demands delicate hyperparameter tuning, which may incur considerable effort and potential inefficiencies in practical implementations. A feasible alternative is to adopt a relative-type error criterion. The term “relative” refers to controlling the errors in minimizing the subproblems based on certain computable quantities related to the current status of the algorithm. This relative-type criterion eliminates the need for a summable tolerance parameter sequence, thereby enhancing the practical implementability of the algorithm. The adoption of a relative-type error criterion started from the seminal works of Solodov and Svaiter [45, 46, 47, 48, 49] on inexact versions of PPA, and has since influenced the development of inexact versions of numerous algorithms, including the variable metric PPA [33], ALM [14], ADMM [15], FISTA [4], and the Bregman-type methods [56, 59]. However, to the best of our knowledge, research on relative-type inexact versions of the pALM remains in its early stage, with limited exploration in [13, 20, 58]. In particular, all these inexact versions are derived by applying a (variable metric) hybrid proximal extragradient method [33, 45, 46] to a certain primal-dual solution mapping of the convex constrained problem, and therefore require an additional correction step to ensure convergence. Consequently, the algorithms developed in [13, 20, 58] essentially deviate from the vanilla pALM framework (1.4). Moreover, our numerical results also indicate that the correction step tends to slow down the convergence of the pALM and thus increase the computational cost. These observations naturally lead us to raise the following question:
@@
In light of this, the primary goal of this paper is to address the above question by designing a relative-type error criterion for the vanilla pALM framework (1.4). Note that the relative-type error criterion proposed by Eckstein and Silva [14] for the ALM innovatively introduces an auxiliary error variable to control the inexactness while ensuring convergence. In this work, we shall demonstrate that a similar idea can be adapted to the vanilla pALM, but with essential modifications. We should emphasize that the corresponding theoretical convergence analysis also requires substantial modifications, due to the inclusion of the proximal term. Interestingly, while the initial motivation for considering pALM was to ensure the nonsingularity of the (generalized) Hessian of the ALM subproblem’s objective function at its solution point, our analysis reveals that this proximal term can further improve the existing convergence properties of the relative-type inexact ALM studied in [1, 14, 61]. Specifically, under standard assumptions, it ensures the convergence of both primal and dual sequences, as well as their asymptotic (super)linear convergence rates. The key contributions and findings of this paper are summarized as follows:
@@
  • We develop a relative-type inexact proximal augmented Lagrangian method (ripALM) to solve the dual problem (1.2). To the best of our knowledge, the proposed ripALM is the first inexact version of the vanilla pALM (1.4). It not only shares the same convergence properties for the sequence $\{\bm{x}^{k}\}$ as the relative-type inexact ALM studied in [1, 14, 61], but also offers additional theoretical advantages: the sequence $\{\bm{y}^{k}\}$ is guaranteed to be bounded; moreover, under a relatively weaker [^2] As noted in [24, Lemma 2.4], the error bound condition in Assumption A used in this work is relatively weaker than the local upper Lipschitz continuity of $(\partial\ell)^{-1}$ at the origin. The latter condition was used in [61] to establish the asymptotic (super)linear convergence rate for the relative-type inexact ALM. For this reason, we use the term “weaker” here. error bound assumption, the whole sequence $\{\bm{y}^{k}\}$ is further proved to be convergent, and both primal and dual sequences achieve asymptotic (super)linear convergence rates. These results provide stronger theoretical support for the applicability of ripALM, particularly in scenarios where accurate solutions of both primal and dual problems are required.
@@
  • Numerically, we implement the proposed ripALM and apply it to solve quadratically regularized optimal transport (QROT) problems and basis pursuit denoising (BPDN) problems. Extensive numerical results validate the promising performance of the proposed ripALM for solving large-scale QROT and BPDN problems. Moreover, comparisons with existing methods underscore our motivation and contributions in developing a relative-type error criterion for the vanilla pALM. In addition, since problem (1.1) has strong modeling capability, our ripALM can potentially be extended to solve optimal transport problems with other convex regularizers such as the group quadratic regularizer [58], provided that the regularizers are proximal-friendly and the generalized Jacobians associated with the pALM-subproblems are easy to compute. Given the growing interests in developing efficient algorithms for computational optimal transport [34], our work may open up new possibilities for applying the powerful pALM algorithmic framework to a wide variety of real-world applications.
@@
The remaining parts of this paper are organized as follows. Section 2 describes the main algorithmic framework of ripALM, whose convergence analysis is conducted in Section 3. Section 4 showcases how to apply the proposed ripALM for solving QROT and BPDN problems. Numerical experiments are conducted in Section 5. Finally, some concluding remarks are summarized in Section 6.
@@
Notation. We use $\mathbb{R}^{n}$, $\mathbb{R}_{+}^{n}$, $\mathbb{R}^{m\times n}$ and $\mathbb{R}_{+}^{m\times n}$ to denote the sets of $n$-dimensional real vectors, $n$-dimensional non-negative vectors, $m\times n$ real matrices and $m\times n$ real non-negative matrices, respectively. For a vector $\bm{x}\in\mathbb{R}^{n}$, $x_{i}$ denotes its $i$-th entry, $\|\bm{x}\|$ denotes its Euclidean norm, and $\|\bm{x}\|_{H}:=\sqrt{\langle\bm{x},\,H\bm{x}\rangle}$ denotes its weighted norm associated with a symmetric positive definite matrix $H\in\mathbb{R}^{n\times n}$. For a matrix $X\in\mathbb{R}^{m\times n}$, $X_{ij}$} denotes its $(i,j)$-th entry, $\|X\|_{F}$ denotes its Frobenius norm, and $\texttt{vec}(X)$ denotes the vectorization of $X$, where $\texttt{vec}(X)]_{i+(j-1)m}=X_{ij}$ for any $1\leq i\leq m$ and $1\leq j\leq n$. For simplicity, given an integer $n>0$, we use $\bm{1}_{n}\in\mathbb{R}^{n}$ to denote the $n$-dimensional vector of all ones, and use $I_{n}$ to denote the $n\times n$ identity matrix.
@@ 2 A relative-type inexact pALM
In this section, we focus on developing a relative-type inexact proximal augmented Lagrangian method (ripALM) to solve the dual problem (1.2). The algorithmic framework is developed based on the parametric convex duality framework (see, for example, [39, 40] and [43, Chapter 11]).
@@
Note that $G$ is proper closed convex since $f^{*}$ is proper closed convex. Then, the (ordinary) Lagrangian function of problem (1.2) can be defined by taking the concave conjugate of $G$ with respect to its second argument (see [43, Definition 11.45]), that is,
@@
Clearly, $\ell$ is convex in its first argument and concave in the second argument. For a given penalty parameter $\sigma>0$, the augmented Lagrangian function of problem (1.2) is defined by (see [43, Example 11.57])
$\mathcal{L}_{\sigma}(\bm{y},\,\bm{x})$= $\sup\limits_{\bm{s}\in\mathbb{R}^{N}}\left\{\ell(\bm{y},\bm{s})-\frac{1}{2\sigma}\|\bm{s}-\bm{x}\|^{2}\right\}$
@@
Here, we assume for the remainder of this paper that $\mathcal{L}_{\sigma}$ denotes the augmented Lagrangian function of the dual problem (1.2), omitting the superscript “dual” as in (1.3).
@@
With the above preparations, we are now ready to present our ripALM for solving problem (1.2) in Algorithm 1.
@@
Input: $\rho\in[0,1)$, $\{\sigma_{k}\}_{k=0}^{\infty}$ such that $\inf_{k\geq 0}\{\sigma_{k}\}>0$, and $\{\tau_{k}\}_{k=0}^{\infty}$ such that $0<\inf_{k\geq 0}\{\tau_{k}\}\leq\sup_{k\geq 0}\{\tau_{k}\}<\infty$. Choose $\bm{y}^{0},\,\bm{w}^{0}\in\mathbb{R}^{M}$ and $\bm{x}^{0}\in\mathbb{R}^{N}$ arbitrarily. Set $k=0$.
@@
    $\leq$ $\;\rho\left(\big\|\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}{A}^{\top}\bm{y}^{k+1}\big)-\bm{x}^{k}\big\|^{2}+\tau_{k}\big\|\bm{y}^{k+1}-\bm{y}^{k}\big\|^{2}\right),$
@@
    $\bm{x}^{k+1}=\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}{A}^{\top}\bm{y}^{k+1}\big),\quad\bm{w}^{k+1}=\bm{w}^{k}-\sigma_{k}\Delta^{k+1}.$ (2.6)
@@
In line with pALM-type methods, at each iteration, our ripALM in Algorithm 1 approximately minimizes the sum of the augmented Lagrangian function $\mathcal{L}_{\sigma_{k}}(\cdot,\,\bm{x}^{k})$ and a proximal term $\frac{\tau_{k}}{2\sigma_{k}}\big\|\cdot-\bm{y}^{k}\big\|^{2}$ under the error criterion (2.5), followed by the updates of the multiplier $\bm{x}^{k}$ and the auxiliary error variable $\bm{w}^{k}$. Next, we show that the error criterion (2.5) is achievable at each iteration, provided that a convergent subsolver is employed. Specifically, the subproblem (2.4) admits a unique solution $\bm{y}^{k,*}$ due to the strong convexity, which satisfies
$\nabla_{y}\mathcal{L}_{\sigma_{k}}(\bm{y}^{k,*},\bm{x}^{k})+\tau_{k}\sigma_{k}^{-1}(\bm{y}^{k,*}-\bm{y}^{k})=\bm{0}.$
Then, a suitable subsolver (e.g., a gradient-type or Newton-type method) for solving the subproblem (2.4) can generate a sequence $\{\bm{y}^{k,t}\}$ converging to $\bm{y}^{k,*}$ as $t\to\infty$. Since the objective function in (2.4) is continuously differentiable (by the property of the Moreau envelope [43]), it follows that
$\Delta^{k,t}:=\nabla_{y}\mathcal{L}_{\sigma_{k}}(\bm{y}^{k,t},\bm{x}^{k})+\tau_{k}\sigma_{k}^{-1}(\bm{y}^{k,t}-\bm{y}^{k})\to\bm{0}\quad\mathrm{as}\quad t\to\infty,$
which further implies that
$2\big|\langle\bm{w}^{k}-\bm{y}^{k,t},\,\sigma_{k}\Delta^{k,t}\rangle\big|+\big\|\sigma_{k}\Delta^{k,t}\big\|^{2}\to 0\quad\mathrm{as}\quad t\to\infty.$
Therefore, the left-hand side of (2.5) converges to 0 as $t\to\infty$. On the other hand, if $\bm{y}^{k}$ is not the unique solution of the subproblem (2.4) with $\tau_{k}\neq 0$ (which is the case of interest, as otherwise we can directly proceed to the next iteration), then the term $\tau_{k}\|\bm{y}^{k,t}-\bm{y}^{k}\|^{2}$ and thus the right-hand side of (2.5) would not approach 0 as $t\to\infty$. Consequently, in this case, (2.5) must be satisfied at some iterate $\bm{y}^{k,t}$ after finitely many $t$ inner iterations. Once (2.5) is satisfied at $\bm{y}^{k,t}$, we set $\bm{y}^{k+1}:=\bm{y}^{k,t}$ and proceed to the next outer iteration.
@@
Compared to the recent semismooth Newton based inexact proximal augmented Lagrangian (Snipal) method developed in [24, Section 3], our ripALM in Algorithm 1 employs a significantly different error criterion (2.5) for solving the subproblem (2.4). Specifically, in our context, Snipal requires the error term $\Delta^{k+1}$ to satisfy
$(A)~~\big\|\Delta^{k+1}\big\|\leq\frac{\min\left(\sqrt{\tau_{k}},\,1\right)}{\sigma_{k}}\varepsilon_{k},~~\varepsilon_{k}\geq 0,~~\sum_{k=0}^{\infty}\varepsilon_{k}<\infty,$ (2.7)
$(B)~~\big\|\Delta^{k+1}\big\|\leq\frac{\delta_{k}\min\left(\sqrt{\tau_{k}},\,1\right)}{\sigma_{k}}\sqrt{\big\|\Delta_{x}^{k+1}\big\|^{2}+\tau_{k}\big\|\Delta_{y}^{k+1}\big\|^{2}},~~0\leq\delta_{k}<1,~~\sum_{k=0}^{\infty}\delta_{k}<\infty,$